OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#5909 — Lille loop 2 (HEL59,TRI59, ASQ59, LZI59,LQN59)- 20/10/11
Scheduled Maintenance Report for Web Cloud
Completed
We will intervene on the loop 2 for security works (same as what has been done on the Loop 1, see task # 5763). This involves a migration of DSLAMs uplinks to a new transportation equipment. So there will be a cut off for the service on each successive NRA,we will try to minimize it as much as we can.
We planned a maintenance window of 1 hour by NRA, knowing that the effective cut off should not take more than few minutes. Work will begin around 22:00. The NRAs will be impacted in the following order: HEL59, TRI59, ASQ59, LZI59, LQN59.

Update(s):

Date: 2011-10-21 09:01:14 UTC
LQN switched.

Date: 2011-10-21 09:00:54 UTC
LZI59 switched. We prepare for the switch of LQN59.

Date: 2011-10-21 09:00:18 UTC
We will switch LZI59.

Date: 2011-10-21 09:00:02 UTC
ASC59 switched.

Date: 2011-10-21 08:59:38 UTC
We will switch ASC59.


Date: 2011-10-21 08:57:26 UTC
We will continue the maintenance tasks tonight starting from 22:00. Expected switch of NRAs ASC, LZI and LQN.


Date: 2011-10-19 03:31:32 UTC
Our optical equipment is up on ASC but the link is not prodable, there is an amp to settle. We will not switch ASC tonight.
Works completed. We will continue in a few days with some failovers for ASC, LZI and LQN.

Date: 2011-10-19 03:28:06 UTC
TRI59 switched,we are preparing the link TRI<>ASC.

Date: 2011-10-19 03:27:39 UTC
HEL59 switched,we are preparing the switching of TRI59.

Date: 2011-10-19 03:26:51 UTC
We are switching HEL59.

Date: 2011-10-19 03:26:26 UTC
Works are ongoing, we are preparing the switchnig on HEL59.

Date: 2011-10-19 03:25:28 UTC
Securing works begin tonight at 22:00. We will migrate the following sites: HEL59, TRI59 and ASC59.
Posted Oct 19, 2011 - 03:24 UTC